A production process sampled 24 times with a sample size of 6 yields an overall mean of 25.2 and an average range of 0.8. Calculate the lower control limit for the mean based on 3 standard deviations.

a. 24.8

b. Cannot be determined with the information given

c. 25.6

d. 25.1

e. 22.8
Answer #1

for sample size of 6

contro chart constant for X-chart = A2 = 0.483

X-bar-bar = 25.2

R-bar = 0.8

X bar chart lower control limit =

LCL(X-bar)= x-bar-bar - (a2 * R-bar)

= 25.2 - (0.483*0.8)

= 24.81

option - A